# Distributed Primal-Dual Algorithms: Unification, Connections, and Insights

### v1

@@ 5.1 Experiment 1
(b) Proximal-1-PD, fixed $\eta_{1}$
(c) Proximal-2-PD, fixed $\rho$
(d) Proximal-2-PD, fixed $\eta_{2}$
(e) LinConsensus-PD, fixed $\beta$
@@
5.2 Experiment 2
In this experiments, we test five the update rules on Ridge Regression problem and LASSO problem, where each $\ell_{i}=\frac{1}{2}(y_{i}-x_{i}^{\top}w)^{2}$, using synthetic data.
Dataset. We generate $n=3000$ training examples $\{x_{i},y_{i}\}_{i=1}^{n}$ according to the model
$y_{i}=\langle x_{i},w^{*}\rangle+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,1),$
where $x_{i}\in\mathbb{R}^{d}$ with $d=500$. The samples are distributed uniformly on $K=30$ machines. We set $w^{*}$ as the vector of all ones, whereas for the $\ell_{1}$ penalty, we let the first 100 elements of $w^{*}$ be ones and the rest be zeros.
For the generation of $x_{i}$’s, we designed two cases: IID data and non-IID data. (1) Under the IID setting, we generate each $x_{i}\sim\mathcal{N}(0,\Sigma)$ where the covariance matrix $\Sigma$ is diagonal with $\Sigma_{j,j}=j^{-2}$. This covariance setting renders an ill-conditioned dataset, making it a challenging situation of solving distributed and large-scale optimization problem. (2) To generate the non-IID case, we follow a setup similar to the one in (Zhou & Li, 2023). Specifically, we generate $[n/3]$ samples $x_{i}$ from the standard normal distribution, $[n/3]$ samples from the Student’s $t$-distribution with 5 degrees of freedom, and the rest samples are from the uniform distribution on $[-5,5]$. After generating all the samples, we shuffle them and randomly distribute them across $K$ machines.
Method. We run the five update rules to solve the Ridge Regression problem on IID and non-IID dataset, and run the four ADMM algorithms to solve the LASSO problem. In these update rules, we select the suggested value of $\gamma,\sigma,\eta_{1},\eta_{2},$ and $\tau$, and select the optimal $\beta$ or $\rho$ to achieve optimal performance. Notably, it has been observed that the optimal $\beta$ and $\rho$ in Consensus-PD and Proximal-1-PD satisfies $\beta K=\rho^{-1}$, and so are the optimal $\beta$ and $\rho$ in LinConsensus-PD and Proximal-2-PD, indicating their connection.
@@ 5.2 Experiment 2
5.3 Binary Classification with Real Data
Datasets.
The real datasets from the LibSVM library (Chang & Lin, 2011) used in the study are a1a, w8a, and real-sim. Details of each dataset, including the number of samples and features are summarized in Table 1. In our experiments, all samples in each dataset are evenly distributed on the machines. We select different numbers of machines for each dataset to evaluate the performance of the proposed approach, where the number of machines is also given in Table 1.
Method.
Results.
(a) Data: a1a, SVM+Ridge
(b) Data: a1a, SVM+$\ell_{1}$
(c) w8a, SVM+Ridge
(d) w8a, SVM+$\ell_{1}$
(e) real-sim, SVM+Ridge
(f) real-sim, SVM+$\ell_{1}$
Figure 4: Relative gap differences versus the number of communication rounds for various real datasets across different models. The first row of plots illustrates the results for SVM with a lasso penalty across different datasets, while the second row shows the results for SVM with a ridge penalty across the same datasets.
Table 1: Description of the datasets.
Dataset $n$ $d$ $K$
a1a 1605 119 10
w8a 49749 300 60
real-sim 72309 20958 100
@@ 6 Conclusion
In this article, we unified distributed primal-dual algorithms, including CoCoA, two proximal ADMM algorithms, consensus ADMM, and linearized ADMM into updates rule that only involve the primal and dual variable updates. Among them, the two proximal ADMM algorithms are new, obtained from choosing two positive definite matrices to enable the proximal ADMM algorithm to solve distributed, regularized federated learning problem. The unified update rules reveal that the CoCoA algorithm can be interpreted as a special case of proximal ADMM with a specific tuning parameter, and proximal ADMM and consensus ADMM are strongly related and have similar convergence property. This framework enables the use of the gap between the primal and dual objectives as a stopping criterion for the consensus ADMM algorithm. The framework also enables us to use a simple and unified ergodic convergence analysis for ADMM variants. By thoroughly investigating the influence of tuning parameters on convergence speed, we found that all ADMM variants consistently outperform the CoCoA-PD with properly selected tuning parameters, as validated by the experiments with synthetic and real-world datasets.
One missing stone in the connection between the algorithms we investigated is an accurate characterization of the connection between Consensus-PD and Proximal-1-PD, and between LinConsensus-PD and Proximal-2-PD, as both experiments and transformation have indicated strong associations between both pairs. Another interesting task is to investigate and include other primal dual algorithms into this category, identify their relationships and and describe their common characteristics. We will leave them for future studies.
@@ Appendix A Derivation of the Dual Problem
Let $w^{\top}x_{i}=u_{i}$ for any $i=1,\ldots,n$, we can equivalently transform the original problem (1) as the following form:
@@
[…]
@@
For the distributed problem form (2), the corresponding distributed dual problem form is thus
@@ Appendix B Proofs for the Results in Section 2
$\underset{\tilde{v}_{[k]}\in\mathbb{R}^{n_{k}}}{\min}\;\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(\tilde{v}_{i})+\frac{1}{2n^{2}\beta}\left(\tilde{v}_{[k]}-\tilde{v}_{[k]}^{(t)}\right)^{\top}X_{[k]}^{\top}X_{[k]}\left(\tilde{v}_{[k]}-\tilde{v}_{[k]}^{(t)}\right)-\frac{1}{n}\Big{\langle}X_{[k]}^{\top}\left(w^{(t)}+\frac{1}{\beta}u_{k}^{(t)}-\frac{1}{n\beta}X_{[k]}\tilde{v}_{[k]}^{(t)}\right),\tilde{v}_{[k]}-\tilde{v}_{[k]}^{(t)}\Big{\rangle}.$ (13)
@@
To complete the proof, we need to show that the dual variable $\tilde{v}$ converges to the global dual variable $v$. The details of this convergence will be addressed in the convergence theory section. By further linearizing the local data matrix $X_{[k]}^{\top}X_{[k]}$ in the dual variable $v$ update, we derive the corresponding update formula for the consensus ADMM incorporating linearization techniques. ∎
B.2 Proof of Proposition 2.2
@@
Appendix C Proofs for the Results in Section 3
C.1 Proof of Corollary 3.1
@@ Appendix C Proofs for the Results in Section 3
Compared to the CoCoA-PD update, this update formula matches it when $\eta_{1}=\sigma$, but differs in the $w$-update.
C.2 Proof of Proposition 3.2
@@
[…]
@@
00 0 $\in$ $\partial g(w^{(t+1)})+\rho^{-1}\left(w^{(t+1)}-w^{(t)}+\frac{\rho}{n}Xv^{(t)}\right).$
00 0 $\in$ $\frac{1}{n}\partial\ell^{*}(v^{(t+1)})+\frac{\rho\eta_{1}}{n^{2}}\mbox{diag}\left(X_{[1]}^{\top}X_{[1]},\ldots,X_{[K]}^{\top}X_{[K]}\right)(v^{(t+1)}-v^{(t)})-\frac{1}{n}X^{\top}\left(2w^{(t+1)}-w^{(t)}\right),$
@@
[…]
@@
For the standard consensus ADMM, we could equivalently write the updates in the Proposition 2.1 by utilizing the first-order optimality conditions as follows:
00 0 $\in$ $\partial g(w^{(t+1)})+\beta K\left(w^{(t+1)}-w^{(t)}+\frac{1}{n\beta K}X\left(2v^{(t+1)}-v^{(t)}\right)\right),$
00 0 $\in$ $\frac{1}{n}\partial\ell^{*}(v^{(t+1)})+\frac{1}{n^{2}\beta}\mbox{diag}\left(X_{[1]}^{\top}X_{[1]},\ldots,X_{[K]}^{\top}X_{[K]}\right)(v^{(t+1)}-v^{(t)})-\frac{1}{n}X^{\top}w^{(t)}.$
@@
Appendix D Proofs for the Results in Section 4
D.1 Proof of Theorem 4.1
@@ Appendix D Proofs for the Results in Section 4
$L(w^{(t+1)};v)-L(w;v^{(t+1)})$ $=$ $L(w^{(t+1)};v)-L(w^{(t+1)};v^{(t+1)})+L(w^{(t+1)};v^{(t+1)})-L(w;v^{(t+1)})$
$\leq$ $\langle u^{(t+1)},z^{(t+1)}-z\rangle=\left(z^{(t)}-z^{(t+1)}\right)^{\top}P\left(z^{(t+1)}-z\right)$
$=$ $\frac{1}{2}\|z^{(t)}-z\|_{P}^{2}-\frac{1}{2}\|z^{(t+1)}-z\|_{P}^{2}-\frac{1}{2}\|z^{(t)}-z^{(t+1)}\|_{P}^{2}$
$\leq$ $\frac{1}{2}\|z^{(t)}-z\|_{P}^{2}-\frac{1}{2}\|z^{(t+1)}-z\|_{P}^{2},$
@@
We first show that $\{z^{(t)}\}$ is bounded.
Let $z^{*}=(w^{*},v^{*})$ denote the optimal solution of the saddle point problem (8). Firstly, from the convexity-concavity property of the objective function $L(w;v)$, we have that
$L(w^{(t+1)};v)-L(w;v^{(t+1)})$ $=L(w^{(t+1)};v)-L(w^{(t+1)};v^{(t+1)})+L(w^{(t+1)};v^{(t+1)})-L(w;v^{(t+1)})$ (14)
@@
$L(\bar{w}^{(T)};v^{*})-L(w^{*};\bar{v}^{(T)})\leq\frac{1}{T}\sum_{t=0}^{T-1}\left\{L(w^{(t+1)};v^{*})-L(w^{*};v^{(t+1)})\right\}\leq\frac{1}{2T}\|z^{(0)}-z^{*}\|^{2}_{P}+\frac{D\sum_{t=1}^{T}\|\epsilon^{(t)}\|}{T},$

### v2

@@ 6 Experiments
(b) Proximal-2-PD, fixed $\rho$
(c) LinConsensus-PD, fixed $\beta$
(d) Proximal-1-PD, fixed $\eta_{1}$
(e) Proximal-2-PD, fixed $\eta_{2}$
@@
Experiment 2. We test five the update rules on Ridge Regression problem and LASSO problem, where each $\ell_{i}=\frac{1}{2}(y_{i}-x_{i}^{\top}w)^{2}$, using synthetic data. The data generation mechanism is detailed in Appendix D.1. We run the five update rules to solve the Ridge Regression problem on IID and non-IID dataset, and run the four ADMM algorithms to solve the LASSO problem. In these update rules, we select the suggested value of $\gamma,\sigma,\eta_{1},\eta_{2},$ and $\tau$, and select the optimal $\beta$ or $\rho$ to achieve optimal performance. Notably, it has been observed that the optimal $\beta$ and $\rho$ in Consensus-PD and Proximal-1-PD satisfies $\beta K=\rho^{-1}$, and so are the optimal $\beta$ and $\rho$ in LinConsensus-PD and Proximal-2-PD, indicating their connection.
@@ 7 Conclusion
In this article, we unified distributed primal-dual algorithms, including CoCoA, two proximal ADMM algorithms, consensus ADMM, and linearized ADMM into updates rule that only involve the primal and dual variable updates. Among them, the two proximal ADMM algorithms are new, obtained from choosing two positive definite matrices to enable the proximal ADMM algorithm to solve distributed, regularized federated learning problem. The unified update rules reveal that the CoCoA algorithm can be interpreted as a special case of proximal ADMM with a specific tuning parameter, and proximal ADMM and consensus ADMM are equivalent. The findings in the paper also indicated rich expressiveness of distributed learning that involves global primal updates and local dual updates. This framework enables the use of the gap between the primal and dual objectives as a stopping criterion for the consensus ADMM algorithm, and also enables us to use a simple and unified ergodic convergence analysis for ADMM variants. By thoroughly investigating the influence of tuning parameters on convergence speed, we found that all ADMM variants consistently outperform the CoCoA-PD with properly selected tuning parameters, as validated by the experiments with synthetic and real-world datasets.
Appendix
@@ Appendix A Derivation of the Dual Problem
Let $w^{\top}x_{i}=u_{i}$ for any $i=1,\ldots,n$, we can equivalently transform the original problem (P) as the following form:
@@
$\underset{w,u}{\inf}\;L(w,u;v)$ $=$ $\underset{u}{\inf}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(\ell_{i}(u_{i})-v_{i}u_{i}\right)\right\}+\underset{w}{\inf}\left\{g(w)+\langle w,\frac{1}{n}\sum_{i=1}^{n}v_{i}x_{i}\rangle\right\}$
@@
For the distributed problem form (D), the corresponding distributed dual problem form is thus
@@
Furthermore, the KKT conditions are listed as follows:
$\left\{\begin{aligned} &x_{i}^{\top}w^{*}=u_{i}^{*},\quad i=1,\ldots,n,\\ &v_{i}^{*}\in\partial\ell_{i}(u_{i}^{*}),\quad i=1,\ldots,n,\\ &-\frac{1}{n}\sum_{i=1}^{n}v_{i}^{*}x_{i}\in\partial g(w^{*}).\end{aligned}\right$.
After simplification, we have
$\left\{\begin{aligned} x_{i}^{\top}w^{*}&=\text{Prox}_{\ell_{i}}\left(x^{\top}_{i}w^{*}+v_{i}^{*}\right),\,\quad\mbox{ for any }i=1,\ldots,n,\\ w^{*}&=\text{Prox}_{g}\left(w^{*}-\frac{1}{n}\sum_{i=1}^{n}v_{i}^{*}x_{i}\right).\end{aligned}\right$.
@@ B.1 Proof of Proposition 1
$\begin{split}\underset{\tilde{v}_{[k]}\in\mathbb{R}^{n_{k}}}{\min}\;&\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(\tilde{v}_{i})+\frac{1}{2n^{2}\beta}\left(\tilde{v}_{[k]}-\tilde{v}_{[k]}^{(t)}\right)^{\top}X_{[k]}^{\top}X_{[k]}\left(\tilde{v}_{[k]}-\tilde{v}_{[k]}^{(t)}\right)\\ &-\frac{1}{n}\Big\langle X_{[k]}^{\top}\left(w^{(t)}+\frac{1}{\beta}u_{k}^{(t)}-\frac{1}{n\beta}X_{[k]}\tilde{v}_{[k]}^{(t)}\right),\tilde{v}_{[k]}-\tilde{v}_{[k]}^{(t)}\Big\rangle.\end{split}$ (8)
@@
To complete the proof, we need to show that the dual variable $\tilde{v}$ converges to the global dual variable $v$. The details of this convergence will be addressed in the later section. By further linearizing the local data matrix $X_{[k]}^{\top}X_{[k]}$ in the dual variable $v$ update, we derive the corresponding update formula for the consensus ADMM incorporating linearization techniques.
B.2 Proof of Proposition 2
@@ B.3 Proof of Lemma 1
B.4 Proof of Corollary 1
Proof.
@@ B.4 Proof of Corollary 1
Compared to the CoCoA-PD update, this update formula matches it when $\eta_{1}=\sigma$, but differs in the $w$-update. ∎
B.5 Proof of Corollary 2
Proof.
To see the equivalence between Consensus-PD and Proximal-1-PD algorithms, let’s consider both algorithms applied to the following saddle-point formulation of the general empirical risk minimization problem:
$\min_{w}\max_{v}\,L(w,v):=\frac{1}{n}\sum_{i=1}^{n}\left(v_{i}\langle w,x_{i}\rangle-\ell_{i}^{*}(v_{i})\right)+g(w),$
where the loss function $\ell_{i}(w^{\top}x_{i})$ is represented in its convex conjugate form as $\ell_{i}(w^{\top}x_{i})=\underset{v_{i}\in\mathbb{R}}{\sup}\left\{v_{i}\langle w,x_{i}\rangle-\ell_{i}^{*}(v_{i})\right\}$. The Consensus-PD update (see Proposition 1 of our paper) is
$v^{(t+1)}_{[k]}=\underset{v_{[k]}\in\mathbb{R}^{n_{k}}}{\arg\min}\;\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})-\frac{1}{n}\Big\langle X_{[k]}^{\top}w^{(t)},v_{[k]}\Big\rangle+\frac{1}{2n^{2}\beta}\Big\|v_{[k]}-v_{[k]}^{(t)}\Big\|_{X_{[k]}^{\top}X_{[k]}}^{2},\quad k\in[K]$
$w^{(t+1)}=\mathrm{prox}_{(\beta K)^{-1}g}\left(w^{(t)}-\frac{1}{n\beta K}X\left(2v^{(t+1)}-v^{(t)}\right)\right).$
It can be equivalently written as
$\begin{split}v^{(t+1)}&=\underset{v}{\arg\max}\,L(w^{(t)},v)-\frac{s_{1}}{2}\|v-v^{(t)}\|^{2}_{M_{1}},\\ w^{(t+1)}&=\underset{w}{\arg\min}\,L(w,2v^{(t+1)}-v^{(t)})+\frac{s_{2}}{2}\|w-w^{(t)}\|^{2}_{M_{2}},\end{split}$
where $s_{1}=\frac{1}{n^{2}\beta}$, $s_{2}=\beta K$, $M_{1}=\mathrm{diag}\left(X_{[1]}^{\top}X_{[1]},\ldots,X_{[K]}^{\top}X_{[K]}\right)$, and $M_{2}=I$ are the algorithm dependent parameters.
Now, consider instead solving the problem $\underset{v}{\max}\underset{w}{\min}\,L(w,v)$, which is equivalent with solving $\underset{v}{\min}\underset{w}{\max}\,-L(w,v)$. The same algorithm Consensus-PD can be applied on this problem, leading to:
$\begin{split}w^{(t+1)}&=\underset{w}{\arg\max}\;-L(w,v^{(t)})-\frac{s_{2}}{2}\|w-w^{(t)}\|^{2}_{M_{2}},\\ v^{(t+1)}&=\underset{v}{\arg\min}\;-L(2w^{(t+1)}-w^{(t)},v)+\frac{s_{1}}{2}\|v-v^{(t)}\|^{2}_{M_{1}}.\end{split}$
After some algebraic operations, we can equivalently write the above iterative form as
$v^{(t+1)}_{[k]}=\underset{v_{[k]}\in\mathbb{R}^{n_{k}}}{\arg\min}\,\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})-\frac{1}{n}\Big\langle X_{[k]}^{\top}\left(2w^{(t)}-w^{(t-1)}\right),v_{[k]}\Big\rangle+\frac{1}{2n^{2}\beta}\Big\|v_{[k]}-v_{[k]}^{(t)}\Big\|_{X_{[k]}^{\top}X_{[k]}}^{2},$
$w^{(t+1)}=\mathrm{prox}_{(\beta K)^{-1}g}\left(w^{(t)}-\frac{1}{n\beta K}Xv^{(t+1)}\right).$
which is the same as the Proximal-1-PD update form (see Equation (3) of Proposition 2) under the parameter setting $\eta_{1}=\eta^{*}=K$ and $\rho=\frac{1}{\beta K}$.
Because of the convex-concave structure of the saddle-point function $L(w,v)$, we know that $\underset{w}{\min}\underset{v}{\max}\,L(w,v)$ and $\underset{v}{\max}\underset{w}{\min}\,L(w,v)$ have the same solution. Therefore, in summary, Proximal-1-PD is just to use Consensus-PD to solve the max-min problem. A similar derivation holds for LinConsensus-PD and Proximal-2-PD. Thus, we verify the Corollary 2. ∎
Appendix C Proofs for the Results in Section 5
C.1 Proof of Lemma 2
@@
$\quad-\frac{1}{n}\Big\langle X_{[k]}^{\top}\left(2w^{(t+1)}-w^{(t)}\right),v_{[k]}\Big\rangle,\quad k\in[K]\;\text{(in parallel)}.$
@@
$0\in\partial g(w^{(t+1)})+\rho^{-1}\left(w^{(t+1)}-w^{(t)}+\frac{\rho}{n}Xv^{(t)}\right),$
$0\in\frac{1}{n}\partial\ell^{*}(v^{(t+1)})+\frac{\rho\eta_{1}}{n^{2}}\mbox{diag}\left(X_{[1]}^{\top}X_{[1]},\ldots,X_{[K]}^{\top}X_{[K]}\right)(v^{(t+1)}-v^{(t)})-\frac{1}{n}X^{\top}\left(2w^{(t+1)}-w^{(t)}\right).$
@@
$v_{[k]}^{(t+1)}$ $=\mathrm{prox}_{(n/\rho\eta_{2})\ell^{*}_{[k]}}\left(v_{[k]}^{(t)}+\frac{n}{\rho\eta_{2}}X_{[k]}^{\top}\left(2w^{(t+1)}-w^{(t)}\right)\right),\quad k\in[K]\;\text{(in parallel)}.$
@@
For the standard consensus ADMM, we could equivalently write the updates in the Proposition 1 by utilizing the first-order optimality conditions as follows:
$0$ $\in$ $\partial g(w^{(t+1)})+\beta K\left(w^{(t+1)}-w^{(t)}+\frac{1}{n\beta K}X\left(2v^{(t+1)}-v^{(t)}\right)\right),$
$0$ $\in$ $\frac{1}{n}\partial\ell^{*}(v^{(t+1)})+\frac{1}{n^{2}\beta}\mbox{diag}\left(X_{[1]}^{\top}X_{[1]},\ldots,X_{[K]}^{\top}X_{[K]}\right)(v^{(t+1)}-v^{(t)})-\frac{1}{n}X^{\top}w^{(t)}.$
@@
C.2 Proof of Theorem 1
@@
$\begin{split}&L(w^{(t+1)};v)-L(w;v^{(t+1)})\\ =&L(w^{(t+1)};v)-L(w^{(t+1)};v^{(t+1)})+L(w^{(t+1)};v^{(t+1)})-L(w;v^{(t+1)})\\ \leq&\langle u^{(t+1)},z^{(t+1)}-z\rangle=\left(z^{(t)}-z^{(t+1)}\right)^{\top}P\left(z^{(t+1)}-z\right)\\ =&\frac{1}{2}\|z^{(t)}-z\|_{P}^{2}-\frac{1}{2}\|z^{(t+1)}-z\|_{P}^{2}-\frac{1}{2}\|z^{(t)}-z^{(t+1)}\|_{P}^{2}\\ \leq&\frac{1}{2}\|z^{(t)}-z\|_{P}^{2}-\frac{1}{2}\|z^{(t+1)}-z\|_{P}^{2},\end{split}$
@@ C.3 Proof of Theorem 2
We first show that $\{z^{(t)}\}$ is bounded. Let $z^{*}=(w^{*},v^{*})$ denote the optimal solution of the saddle point problem (SP). Firstly, from the convexity-concavity property of the objective function $L(w;v)$, we have that
$L(w^{(t+1)};v)-L(w;v^{(t+1)})$ (9)
$=L(w^{(t+1)};v)-L(w^{(t+1)};v^{(t+1)})+L(w^{(t+1)};v^{(t+1)})-L(w;v^{(t+1)})$
@@
$L(\bar{w}^{(T)};v^{*})-L(w^{*};\bar{v}^{(T)})\leq\frac{1}{T}\sum_{t=0}^{T-1}\left\{L(w^{(t+1)};v^{*})-L(w^{*};v^{(t+1)})\right\}\\ \leq\frac{1}{2T}\|z^{(0)}-z^{*}\|^{2}_{P}+\frac{D\sum_{t=1}^{T}\|\epsilon^{(t)}\|}{T},$ (10)
@@
Appendix D Experimental Details
D.1 Experiment 2 data generation details
We generate $n=3000$ training examples $\{x_{i},y_{i}\}_{i=1}^{n}$ according to the model $y_{i}=\langle x_{i},w^{*}\rangle+\epsilon_{i},\epsilon_{i}\sim\mathcal{N}(0,1),$ where $x_{i}\in\mathbb{R}^{d}$ with $d=500$. The samples are distributed uniformly on $K=30$ machines. We set $w^{*}$ as the vector of all ones, whereas for the $\ell_{1}$ penalty, we let the first 100 elements of $w^{*}$ be ones and the rest be zeros. For the generation of $x_{i}$’s, we designed two cases: IID data and non-IID data. (1) Under the IID setting, we generate each $x_{i}\sim\mathcal{N}(0,\Sigma)$ where the covariance matrix $\Sigma$ is diagonal with $\Sigma_{j,j}=j^{-2}$. This covariance setting renders an ill-conditioned dataset, making it a challenging situation of solving distributed and large-scale optimization problem. (2) To generate the non-IID case, we follow a setup similar to the one in [26]. Specifically, we generate $[n/3]$ samples $x_{i}$ from the standard normal distribution, $[n/3]$ samples from the Student’s $t$-distribution with 5 degrees of freedom, and the rest samples are from the uniform distribution on [5,5][-5,5]. After generating all the samples, we shuffle them and randomly distribute them across $K$ machines.
D.2 Experiment 3: Binary Classification with Real Data
Datasets.
The real datasets from the LibSVM library [2] used in the study are a1a, w8a, and real-sim. Details of each dataset, including the number of samples and features are summarized in Table 1. In our experiments, all samples in each dataset are evenly distributed on the machines. We select different numbers of machines for each dataset to evaluate the performance of the proposed approach, where the number of machines is also given in Table 1.
Method.
Results.
(a) Data: a1a, SVM+Ridge
(b) Data: w8a, SVM+Ridge
(c) Data: real-sim, SVM+Ridge
[…]
Figure 4: Relative gap differences versus the number of communication rounds for various real datasets across different models. The first row of plots illustrates the results for SVM with a ridge penalty across different datasets, while the second row shows the results for SVM with a lasso penalty across the same datasets.